NESE tape hardware issues
प्रभावित
आंशिक आउटेज से 2:10 PM तक 9:21 PM
आंशिक आउटेज से 2:10 PM तक 9:21 PM
- हल हुआ · UTC
This issue is resolved.
From the NESE team:
The NESE tape system is up and running and functional now. The tape robot had a frame corruption, parts were replaced and it needed to re-discover hardware in each frame which takes a long time. Also there was an issue where most of the fiber switch ports needed to be manually enabled due work that was performed on the robot. Sorry for this extended downtime.
- पहचाना गया · UTC
The NESE tape robot and tape system is currently not functional. IBM is onsite tech (Saturday). Parts are required and ordered, but no immediate ETA.
The Spectrum Scale disk cache is operational. Data can be written to disk cache, but no data can be sent to tape or retrieved from tape at this time.
Updates as we have them.

